OpenStudy (anonymous):

if cos(t) = 2/3 and t is in the fourth quadrant, then evaluate sin(6pi + t).

OpenStudy (anonymous):

should i square cos and find its sin value and then use the sin(pi + t) = -sin t? :o

OpenStudy (anonymous):

6pi is 3 complete rotations back to the 4th quadrant. We know that sine is negative in that quadrant. From the 2/3 we know that adj/hyp = 2/3. So we can use pythagoreans: \[3^2-2^2=opp^2\] \[9-4=5\] \[opp=\sqrt{5}\] From the definition of sine and knowing that sine is negative we have: \[sin\theta=-\frac{\sqrt{5}}{3}\]

OpenStudy (anonymous):

Clear as mud?

OpenStudy (anonymous):

yes it makes sense. thank you : )

OpenStudy (anonymous):

Happy to help. Keep the signs clear in your head when approaching these problems and view it from a reference angle perspective.

OpenStudy (anonymous):

Okay : D